# Towards Optimal Aggregation of Varying Range Dependencies in Haze Removal

### v2

@@ Abstract
Haze removal aims to restore a clear image from a hazy input. Existing methods have shown significant efficacy by capturing either short-range dependencies for local detail preservation or long-range dependencies for global context modeling. Given the complementary strengths of both approaches, a intuitive advancement is to explicitly integrate them into a unified framework. However, this potential remains underexplored in current research. In this paper, we propose DehazeMatic, which leverages the proposed Transformer-Mamba Dual Aggregation block to simultaneously and explicitly captures both short- and long-range dependencies through dual-path design for improved restoration. To ensure that dependencies at varying ranges contribute optimally to performance, we conduct extensive experiments to identify key influencing factors and determine that an effective aggregation mechanism should be guided by the joint consideration of haze density and semantic information. Building on these insights, we introduce the CLIP-enhanced Dual-path Aggregator, which utilizes the rich semantic priors encapsulated in CLIP and the estimated haze density map, derived from its powerful generalization ability, to instruct the aggregation process. Extensive experiments demonstrate that DehazeMatic outperforms sort-of-the-art methods across various benchmarks.
Figure 1: Illustration of factors influencing the relative importance between short-range and long-range dependencies. We train a Transformer-based dehazing network with a global receptive field and synthesis hazy images with different densities based on the Atmospheric Scattering Model [49]. (a) The horizontal axis represents the Euclidean distance between the current token and all other tokens, while the vertical axis represents the average attention weights between the current token and all tokens at the corresponding distance. The curve data is derived from all tokens of all images in the test set. (b) The bottom visualizes the attention map between the chosen tokens (indicated by the tails of red arrows) and all other tokens. On the left is the visualization of tokens at the same position corresponding to the same ground truth image, but with different haze densities. On the right is a visualization of tokens with different semantic information on the same homogeneous haze image. More analysis is in supplementary material.
1 Introduction
@@
Single image dehazing aims to recover a clear image from a hazy input, serving as a crucial pre-processing step for high-level vision tasks under hazy conditions, such as object detection [39] and semantic segmentation [53].
Figure 2: Illustration of CLIP [52]’s potential to perceive haze and its density. We use the ViT-B/32 OpenCLIP [26] model, pre-trained on the LAION dataset, for our experiments. The values above the images indicate the PSNR and SSIM metrics, which reflect haze density from a quantitative perspective. The values below the images represent the similarity scores between the images and the paired prompts from CLIP. Notably, as haze density increases, the similarity score with the haze-describing prompt also increases.
@@
Current data-driven methods can be categorized into two classes based on the receptive field size of the operators used for feature extraction, namely (i) using convolution [5, 7, 13, 37, 55, 56, 72] or window-based self-attention [35, 34, 59, 64] and (ii) using linear self-attention [51] or State Space Model (SSM) [58, 80, 73]. The former excels at capturing short-range dependencies but lacks effective global modeling capabilities [30, 62, 12]. In contrast, the latter directly models long-range dependencies while ensuring efficient computation; however, it fails to preserve local 2D inductive bias [25, 14].
@@
Both categories of methods have achieved excellent performance, and their respective strengths can address the limitations of the other. Therefore, a natural idea is that explicitly combining short- and long-range dependencies (i.e., leveraging the strengths of both approaches) is expected to yield even more promising results. However, this integration has not been well explored. To address this gap, we will employ a dual-path design that concurrently captures fine-grained perception from nearest neighbor features and holistic contextual cues from non-local relationships. The ensuing problem is that reasonably integrating short- and long-range dependencies from these two paths is a non-trivial task, because tokens with different characteristics in the image require varying ranges of dependencies. Simple addition or concatenation often leads to suboptimal performance. More importantly, there is still no clear guidance on how to determine the relative importance between short-range and long-range dependencies for each token.
One preliminary hypothesis is that haze density serves as a determining factor. To verify this, we conduct an experiment. As shown in Figure 1(a), higher haze density increases reliance on long-range dependencies, whereas lower haze density requires more short-range dependencies. Furthermore, additional investigation reveals that tokens with varying semantic information require different ranges of dependencies, as illustrated on the right side of Figure 1(b). Based on these findings, we conclude that haze density and semantic information jointly determine the relative importance of short-range and long-range dependencies.
In this paper, building on the above insights, we propose DehazeMatic, where Mamba and Transformer are Integrated by CLIP for image dehazing. The core building block of the DehazeMatic is the proposed Transformer-Mamba Dual Aggregation (TrambaDA) block, which employs a dual-path design. In one path, we utilize window-based self-attention [45] to capture short-range dependencies. In the other path, inspired by recent advances in SSMs [18, 19, 44, 82, 42], we incorporate the Selective Scan Space State Sequential Model (S6) block [18], which maintains global information indefinitely with linear complexity. We also adopt the four-way scanning strategy to mitigate the “direction-sensitive" problem caused by the non-causal nature of visual data [44].
Next, to ensure that both short- and long-range dependencies contribute optimally to performance, a dual-path aggregator capable of adaptively learning the relative importance of each path is required. As indicated by Figure 1, the aggregator must capture both the haze density and semantic information of each token in the input image. Recent advances in CLIP [52], pre-trained on web-scale datasets, reveal that it not only encapsulates rich semantic priors (e.g., enabling strong performance in zero-shot semantic segmentation [81, 76]), but also has the potential to perceive haze and its density (as shown in Figure 2). Motivated by this, we propose the CLIP-enhanced Dual-path Aggregator (CedA). Specifically, we treat the affinity scores between the the patch-wise image embeddings from the CLIP image encoder and the text embeddings of paired prompts describing clear and hazy conditions from the CLIP text encoder as the estimated haze density map. To find the best prompts for better estimation and avoid labor-intensive prompt engineering, these paired prompts are learned. For semantic information, we reuse the the patch-wise image embeddings to provide high-level and abstract features while avoiding additional computational overhead. To supplement the low-level details, we incorporate the input features from the TrambaDA blocks and align both representations using a bidirectional cross-attention mechanism. Finally, we generate dual-path aggregation weights based on the refined semantic information and the estimated haze density map.
Figure 3: (a) The architecture of our proposed DehazeMatic. (b) The inner structure of Transformer-Mamba Dual Aggregation (TrambaDA) block. The weights generated by CLIP-enhanced Dual-path Aggregator (CedA) are used to aggregate the outputs of dual-path.
Our contributions can be summarized as follows:
  • We propose DehazeMatic, which demonstrates the superiority of explicitly and simultaneously capturing both short- and long-range dependencies while effectively aggregating dependencies across varying ranges. It achieves SOTA performance in various benchmarks.
  • For the first time, we identify the key factors required for reasonably aggregating dependencies across different ranges in image dehazing. Based on these insights, we propose the CedA module, which performs dynamic aggregation according to haze density and semantic characteristics of each token within the image.
@@
  • The proposed CedA module further explore the potential of CLIP in image dehazing and, for the first time, demonstrate its capability in haze density estimation, all without the need for fine-tuning the encoder.
2 Related Work
2.1 Single Image Dehazing
@@
Due to spatially variant transmission map and atmospheric light, single image dehazing is a highly ill-posed problem. Early prior-based methods [23, 16, 31, 60, 83, 6] achieved dehazing by imposing various assumptions to estimate key parameters in the Atmospheric Scattering Model (ASM) [49]. ASM provides strong theoretical support and interpretability for these methods; however, they often fail when images deviate from statistical laws.
@@
With the rapid advancement of deep learning [33], numerous learning-based methods have been proposed, resulting in improved performance. Early methods [7, 37] employ neural networks to estimate key parameters in the ASM and subsequently restore the haze-free images. Later, ASM-independent deep networks [54, 55, 43, 40, 57, 13, 75, 50, 36, 66, 70, 59, 17, 68, 77, 8, 15, 9, 65, 69, 11] directly estimate clear images or haze residuals. DehazeFormer [59] is a representative method that achieves efficient feature extraction through window-based self-attention and several key modifications. However, its inherently limited receptive field constrains its performance potential. To achieve a global receptive field with lower computational overhead, some methods [58, 71] incorporate frequency domain features into image dehazing, offering the additional benefit of more direct learning of high-frequency information. For the same purpose, other researchers [79, 73] introduce the Mamba [18] architecture into image dehazing. However, none of these methods address how to better integrate global semantic features with local texture features.
2.2 CLIP for Low-Level Vision Tasks
@@
Classic vision-language models, such as CLIP [52], aim to learn aligned features in the embedding space from image-text pairs using contrastive learning. Some studies have explored leveraging the rich prior knowledge encapsulated in CLIP to assist with low-level vision tasks.
@@
In All-in-One image restoration, some researchers [47, 1, 27] use degradation embeddings from the CLIP image encoder to implicitly guide the network in making adaptive responses. In monocular depth estimation, recent studies [74, 4, 24] use CLIP to map patches of the input image to specific semantic distance tokens, which are then projected onto a quantified depth bin for estimation. In low-light enhancement, some methods [41, 48] use text-image similarity between the enhanced results and learnable prompt pairs to train the enhancement network.
Figure 4: Illustration of the CLIP-enhanced Dual-path Aggregator generating aggregation weights. To ensure computational efficiency, the TrambaDA blocks share the same aggregation weights at each stage.
@@
Some works also introduce CLIP into image [63] and video [53] dehazing, but they all use text-image similarity between dehazed results and contrastive prompt sets as regularization to guide the restoration process. In contrast, our method further exploits the potential of CLIP by directly incorporating its latent embeddings into the main network to guide the dehazing process.
3 Method
3.1 Overall Architecture
The architecture of DehazeMatic is depicted in Figure 3. Given a hazy image, shallow features are first extracted and then passed through a symmetric encoder-decoder framework. Each stage of encoder and decoder consists of multiple Transformer-Mamba Dual Aggregation (TrambaDA) blocks, along with corresponding downsampling or upsampling operations. We also incorporate skip connections between the encoder and decoder at each level. Finally, the output of the last decoder stage is combined with the input hazy image to produce the haze-free image.
3.2 Transformer-Mamba Dual Aggregation Block
@@
The TrambaDA block’s backbone employs a dual-path design, with one path capturing short-range dependencies and the other capturing long-range dependencies, thereby providing complementary restoration information for image tokens with distinct characteristics.
Capturing Short-range Dependencies.  We construct this path using window-based self-attention [45], which offers stronger fitting capability through dynamic weights. Specifically, given the input feature $F_{\rm{in}}\in\mathbb{R}^{H\times W\times C}$, we first divide $F_{\rm{in}}$ into non-overlapping windows with size of $M\times M$, then obtain flattened feature $F_{\mathrm{in}}^{i}\in\mathbb{R}^{M^{2}\times C}$ of each window $i$. For simplicity, we only consider the single-head attention. Performing self-attention on non-overlapping windows can be formulated as:
$\begin{split}F_{\mathrm{out\_{short}}}^{i}=&\text{Attention}(F_{\mathrm{in}}^{i}W^{Q},F_{\mathrm{in}}^{i}W^{K},F_{\mathrm{in}}^{i}W^{V}),\\ \text{Attention}&(Q,K,V)=\text{Softmax}(\frac{{QK}^{T}}{\sqrt{d_{k}}})V,\\ F_{\mathrm{out\_{short}}}=&\{F_{\mathrm{out\_{short}}}^{1},F_{\mathrm{out\_{short}}}^{2},\cdots,F_{\mathrm{out\_{short}}}^{N}\},\end{split}$ (1)
where $N=HW/M^{2}$; $W^{Q},W^{K},W^{V}\in\mathbb{R}^{C\times d_{k}}$ are projection matrices for queries, keys, and values, respectively.
Capturing Long-range Dependencies.  We build this path using Mamba [18]’s S6 block, which preserves the global receptive field with linear complexity and incorporates hardware-aware algorithms.
Due to the “direction-sensitive” problem caused by the non-causal nature of visual data, directly applying the S6 block on the flattened image will result in suboptimal performance [44]. Following Vmamba [44], we unfold feature map along four directions, generating four independent sequences. Each feature sequence is processed through the S6 block, and all output sequences are merged into a new sequence. This process can be represented as follows:
$\begin{split}F_{\mathrm{in}}^{d}=expand&(F_{\mathrm{in}},d),d=\{1,2,3,4\},\\ \bar{F}&{}_{\mathrm{in}}^{d}=\mathrm{S6}(F_{\mathrm{in}}^{d}),\\ F_{\mathrm{out\_{long}}}=m&erge(\bar{F}_{\mathrm{in}}^{1},\bar{F}_{\mathrm{in}}^{2},\bar{F}_{\mathrm{in}}^{3},\bar{F}_{\mathrm{in}}^{4}),\end{split}$ (2)
where $expand(\cdot)$ and $merge(\cdot)$ correspond to the scan expand and scan merge operations, respectively.
3.3 CLIP-enhanced Dual-path Aggregator
The CLIP-enhanced Dual-path Aggregator (CedA) utilizes the generated pixel-level weights to adaptively aggregate dual-path outputs, ensuring that dependencies across varying ranges contribute optimally to performance:
@@ 3.3 CLIP-enhanced Dual-path Aggregator
$W_{short},W_{long}\in\mathbb{R}^{H\times W}$ are the aggregation weights corresponding to short- and long-range dependencies, respectively, and their generation process is shown in Figure 4.
3.3.1 Estimating Haze Density Map
Define Estimation Pipeline.  Inspired by Figure 2, we leverage CLIP to estimate the haze density map. Specifically, given a hazy image $I_{\mathrm{haze}}\in\mathbb{R}^{H\times W\times 3}$ and a paired prompts describing haze and clear conditions $T=[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ (e.g., [’hazy image’, ’clear image’]), we encode them into latent space using a pretrained CLIP model:
@@ 3.3.1 Estimating Haze Density Map
Table 1: Quantitative comparisons of various methods on various dehazing benchmarks. The best results and the second best results are in bold and underline, respectively.
Methods Publication RESIDE SOTS (Homogeneous Haze) Non-homogeneous Haze Overhead
Outdoor Indoor Average NH-Haze Dense-Haze
$\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$Param MACs
DCP [22] TPAMI-10 19.14 0.861 16.61 0.855 17.88 0.858 10.57 0.520 10.06 0.385 - -
AOD-Net [37] ICCV-17 24.14 0.920 20.51 0.816 22.33 0.868 15.40 0.569 13.14 0.414 1.76K 0.12G
GridDehazeNet [43] ICCV-19 30.86 0.982 32.16 0.984 31.51 0.983 18.33 0.667 14.96 0.533 0.96M 21.55G
FFA-Net [50] AAAI-20 33.57 0.984 36.39 0.989 34.98 0.987 19.87 0.692 16.09 0.503 4.46M 288.86G
DeHamer [20] CVPR-22 35.18 0.986 36.63 0.988 35.91 0.987 20.66 0.684 16.62 0.560 132.45M 59.25G
SRDefog [28] ACCV-22 - - - - - - 20.99 0.610 16.67 0.500 12.56M 24.18M
MAXIM-2S [61] CVPR-22 34.19 0.985 38.11 0.991 36.15 0.988 - - - - 14.10M 216.00G
SGID-PFF [5] TIP-22 30.20 0.975 38.52 0.991 34.36 0.983 - - - - 13.87M 156.67G
PMNet [70] ECCV-22 34.74 0.985 38.41 0.990 36.58 0.988 20.42 0.730 16.79 0.510 18.90M 81.13G
MB-TaylorFormer-B [51] ICCV-23 37.42 0.989 40.71 0.992 39.07 0.991 - - 16.66 0.560 2.68M 38.50G
MITNet [58] MM-23 35.18 0.988 40.23 0.992 37.71 0.990 21.26 0.712 16.97 0.606 2.73M 16.42G
DehazeFormer [59] TIP-23 34.29 0.983 38.46 0.994 36.38 0.989 20.31 0.761 16.66 0.595 4.63M 48.64G
SCANet [21] CVPR-23 - - - - - - 19.52 0.649 15.35 0.508 2.39M 258.63G
DEANet [8] TIP-24 36.03 0.989 40.20 0.993 38.12 0.991 20.84 0.801 16.73 0.602 3.65M 32.23G
UVM-Net [79] Arxiv-24 34.92 0.984 40.17 0.996 37.55 0.990 - - - - 19.25M 173.55G
OKNet [10] AAAI-24 35.45 0.992 37.59 0.994 36.52 0.993 20.29 0.800 16.85 0.620 4.42M 39.54G
DCMPNet [77] CVPR-24 36.56 0.993 42.18 0.996 39.37 0.995 - - - - 18.59M 80.42G
DehazeMatic - 38.21 0.995 41.50 0.996 39.86 0.996 21.47 0.806 17.28 0.629 4.58M 35.50G
Train Learnable Haze/Clear Prompts.  To find the best prompts for improved estimation and avoid time-consuming prompt engineering, we use learnable prompt tokens instead of manually predefined prompts to describe abstract haze and clear conditions.
The training process consists of two stages. In the first stage, we optimize the learnable paired prompts using cross-entropy loss, enabling them to preliminarily distinguish between haze and clear conditions. Given a hazy and clear image $I_{\mathrm{haze}},I_{\mathrm{clear}}\in\mathbb{R}^{H\times W\times 3}$, we define the first-stage loss $\mathcal{L}_{1}$ as:
@@
In the second stage, our goal is to achieve better prediction of haze density. The most straightforward and effective optimization approach is regression. However, there is no ground-truth density map corresponding to the hazy image in the existing dataset. To address this, we construct triplet data $\{I_{\mathrm{haze}},I_{\mathrm{clear}},I_{\mathrm{density}}\}$ based on Atmospheric Scattering Model [49]. The second-stage loss $\mathcal{L}_{2}$ is defined as:

### v3

@@ Abstract
Haze removal aims to restore a clear image from a hazy input. Existing methods achieve notable success by specializing in either short-range dependencies to preserve local details or long-range dependencies to capture global context. Given the complementary strengths of both, a natural progression is to explicitly integrate them within a unified framework and enable their reasonable aggregation. However, this integration remains underexplored. In this paper, we propose DehazeMatic, which simultaneously and explicitly captures both short- and long-range dependencies through a dual-stream design. To optimize the contribution of dependencies at varying ranges, we conduct extensive experiments to identify key influencing factors and find that an effective aggregation mechanism should be guided by the joint consideration of haze density and semantic information. Building on these insights, we introduce the CLIP-enhanced Dual-path Aggregator, which not only enables the generation of fine-grained haze density maps for the first time, but also produces semantic maps within a shared backbone, ultimately leveraging both to instruct the aggregation process. Extensive experiments demonstrate that DehazeMatic outperforms state-of-the-art methods across multiple benchmarks.
Introduction
@@
Image dehazing aims to recover a clear image from a hazy input, serving as a crucial pre-processing step for high-level vision tasks in hazy conditions, such as object detection (Li et al. 2023) and semantic segmentation (Ren et al. 2024).
@@
Current data-driven methods can be categorized into two classes based on the receptive field size of the operators used for feature extraction, namely (i) using convolution (Bai et al. 2022; Cai et al. 2016; Dong et al. 2020; Li et al. 2017; Ren et al. 2018, 2020; Zhang and Patel 2018) or window-based self-attention (Kulkarni, Phutke, and Murala 2022; Kulkarni and Murala 2023; Song et al. 2023; Wang et al. 2023) and (ii) using linear self-attention (Qiu et al. 2023) or State Space Model (SSM) (Shen et al. 2023; Zhou et al. 2024; Zhang et al. 2024a). The former captures short-range dependencies, which provide fine-grained local perception but lack effective global modeling capabilities (Kim et al. 2023; Veit, Wilber, and Belongie 2016; De and Smith 2020). In contrast, the latter captures long-range dependencies with rich contextual information while ensuring efficient computation, yet fails to preserve local 2D inductive biases (Huang et al. 2024; Dosovitskiy 2020).
@@
Both types of methods have achieved excellent performance, and their respective strengths can compensate for each other’s limitations. This naturally motivates the design of a dual-stream network that explicitly integrates short- and long-range dependencies, thereby harnessing the complementary advantages of both approaches and potentially achieving superior performance. However, reasonably aggregating short- and long-range dependencies from these two paths remains a non-trivial task, as tokens with different characteristics in the image require varying ranges of dependencies. Simple operations such as addition, concatenation, or gating mechanisms (Ren et al. 2018; Zhang et al. 2020a; Hou et al. 2025) often lead to suboptimal performance. More importantly, there is still no clear guidance on how to determine the relative importance between short-range and long-range dependencies for each token.
In this paper, we are the first to identify, through both quantitative and qualitative experiments, the key factors that determine the relative importance of short- and long-range dependencies required by each token: haze density and semantic information. Based on these insights, we propose a dual-stream framework for dehazing, termed DehazeMatic, which simultaneously captures both short- and long-range dependencies. Since feature extraction operators are not the primary focus of this work, their selection is relatively flexible. We adopt window-based self-attention (Liu et al. 2021) and the Selective Scan Space State Sequential Model (S6) block (Gu and Dao 2023) as representative implementations.
Figure 1: Illustration of CLIP (Radford et al. 2021)’s potential to perceive haze and its density. We use the ViT-B/32 OpenCLIP (Ilharco et al. 2021) model, pre-trained on the LAION dataset, for our experiments. The values above the images indicate the PSNR and SSIM metrics, which reflect haze density from a quantitative perspective. The values below the images represent the similarity scores between the images and the paired prompts from CLIP. Notably, as haze density increases, the similarity score with the haze-describing prompt also increases.
Next, based on the identified key factors, we aim to generate a haze density map and a semantic information map to guide the aggregation of dependencies across varying ranges. Moreover, to improve model efficiency, we seek to generate both maps using a shared backbone. However, this objective presents significant challenges: not only is there currently no method capable of accurately estimating haze density maps, but extracting two types of information with substantial modality differences from a single backbone remains inherently difficult. Inspired by recent advances in CLIP (Radford et al. 2021), which is pretrained on web-scale datasets and encapsulates rich semantic priors that enable strong performance in zero-shot semantic segmentation (Zhou et al. 2023; Zhang et al. 2024b), we further observe that CLIP has the potential to perceive haze and its density, as demonstrated in Figure 1. Building on this observation, we propose CLIP-enhanced Dual-path Aggregator (CedA). Specifically, we treat the patch-wise image embeddings from the CLIP image encoder as the semantic map. In parallel, we compute affinity scores between these image embeddings and the text embeddings of paired prompts describing clear and hazy conditions—obtained from the CLIP text encoder—and use these scores as the estimated haze density map. To enable accurate estimation while avoiding labor-intensive prompt engineering, we replace manually designed prompts with learnable ones. Training proceeds in a progressive two-stage manner: we first employ cross-entropy loss, followed by regression-based training on a self-constructed triplet dataset. Finally, based on the generated haze density and semantic maps, the dual-path aggregation weights are computed to reasonably and effectively aggregate short- and long-range dependencies.
Our contributions are summarized as follows:
  • We propose DehazeMatic, showing that once effective aggregation is achieved, explicitly and simultaneously modeling short- and long-range dependencies leads to state-of-the-art performance on multiple benchmarks.
  • We are the first to identify the key factors that determine the relative importance of short- and long-range dependencies in image dehazing. Building on this insight, we propose the CedA module, which dynamically aggregates these dependencies without being restricted to any specific feature extraction operator.
@@
  • We further explore the potential of CLIP in haze removal and, for the first time, achieve accurate estimation of haze density maps, without any fine-tuning of the encoder.
Related Work
Single Image Dehazing
@@
Due to spatially variant transmission map and atmospheric light, single image dehazing is a highly ill-posed problem. Early prior-based methods (He et al. 2016; Fattal 2008; Kim, Park, and Eom 2019; Tan 2008; Zhu, Mai, and Shao 2015; Berman, Treibitz, and Avidan 2018) achieved dehazing by imposing various assumptions to estimate key parameters in the Atmospheric Scattering Model (ASM) (Narasimhan and Nayar 2002). ASM provides strong theoretical support and interpretability for these methods; however, they often fail when images deviate from statistical laws.
@@
With the rapid advancement of deep learning (Krizhevsky, Sutskever, and Hinton 2012), various learning-based methods have been proposed, resulting in improved performance. Early methods (Cai et al. 2016; Li et al. 2017) employ neural networks to estimate key parameters in the ASM and subsequently restore the haze-free images. Later, ASM-independent deep networks (Ren et al. 2016, 2018; Liu et al. 2019; Li et al. 2019; Shao et al. 2020; Dong et al. 2020; Zhang et al. 2020b; Qin et al. 2020; Li et al. 2020; Wu et al. 2021; Ye et al. 2022; Song et al. 2023; Feng et al. 2024; Yang et al. 2024b; Zhang, Zhou, and Li 2024; Chen, He, and Lu 2024; Fang et al. 2024b; Cong et al. 2024; Wang et al. 2024b; Yang et al. 2025; Cui et al. 2025) directly estimate clear images or haze residuals. DehazeFormer (Song et al. 2023) is a representative method that achieves efficient feature extraction through window-based self-attention and several key modifications. However, its inherently limited receptive field constrains its performance potential. To achieve a global receptive field with lower computational overhead, some methods (Shen et al. 2023; Yu et al. 2022) incorporate frequency domain features into image dehazing, offering the additional benefit of more direct learning of high-frequency information. For the same purpose, other researchers (Zheng and Wu 2024; Zhang et al. 2024a) introduce the Mamba (Gu and Dao 2023) architecture into dehazing. However, none of these methods address how to better integrate global semantic features with local texture features.
CLIP for Low-Level Vision Tasks
@@
Classic vision-language models like CLIP (Radford et al. 2021), aim to learn aligned features in the embedding space from image-text pairs using contrastive learning. Some studies have explored leveraging the rich prior knowledge encapsulated in CLIP to assist with low-level vision tasks.
@@
In All-in-One image restoration, some researchers (Luo et al. 2023; Ai et al. 2024; Jiang et al. 2025) use degradation embeddings from the CLIP image encoder to implicitly guide the network in making adaptive responses. In monocular depth estimation, recent studies (Zhang et al. 2022; Auty and Mikolajczyk 2023; Hu et al. 2024) use CLIP to map patches of the input image to specific semantic distance tokens, which are then projected onto a quantified depth bin for estimation. In low-light enhancement, some methods (Liang et al. 2023; Morawski et al. 2024) use text-image similarity between the enhanced results and learnable prompt pairs to train the enhancement network.
@@
Some works also introduce CLIP into image (Wang et al. 2024a) and video (Ren et al. 2024) dehazing, but they all use text-image similarity between dehazed results and contrastive prompt sets as regularization to guide the restoration process. In contrast, our method further exploits the potential of CLIP by directly incorporating its latent embeddings into the main network to guide the dehazing process.
Motivational Experiment
Optimal aggregation of dependencies is essential for a dual-stream network to fully leverage short- and long-range cues for dehazing. To this end, we empirically investigate the key factors that govern their relative importance, thereby enabling more reasonable and effective aggregation.
Quantifying the Importance of Dependencies
Dependency denotes the influence exerted by other tokens on the current token (Bengio, Simard, and Frasconi 1994; Hochreiter and Schmidhuber 1997), and its range can be quantified by the Euclidean distance. For the experiment presented in this section, we train a Transformer model with a global receptive field, where the attention weights assigned by other tokens to the current token in the self-attention mechanism are defined as the importance of the corresponding dependencies.
Experimental Design
We begin by hypothesizing potential key influencing factors and then verify their validity through quantitative and qualitative results. To this end, we select image tokens with differing characteristics in terms of the hypothesized factors as the current tokens, and examine whether the dependency importance at a fixed distance varies accordingly. This setup enables us to reflect the relative importance between short-range and long-range dependencies, as the importance of all other tokens with respect to the current token is normalized using L1 normalization.
Experimental Observations
Haze density is intuitively regarded as a key contributing factor. To validate this, we synthesize hazy images with varying density levels using the Atmospheric Scattering Model (Narasimhan and Nayar 2002) and conduct corresponding experiments. As illustrated in Figure 2, as the haze becomes denser, the relative importance of long-range dependencies increases, while that of short-range dependencies decreases—and vice versa. Quantitative results further confirm this observation.
Semantic information is also hypothesized to be an influential factor, as prior work (Huang et al. 2020) indicates that scenes with different levels of complexity require dependencies at varying ranges. To examine this, we conduct experiments on indoor and outdoor images from the RESIDE dataset (Li et al. 2018), which generally exhibit distinct semantic characteristics. As shown in Figure 2, the quantitative results support our hypothesis, while qualitative results further demonstrate that for tokens with different semantic content within the same image, the relative importance of dependencies at different ranges also differs.
Based on these findings, we are the first to propose that in image dehazing, the relative importance of short-range and long-range dependencies is jointly influenced by both haze density and semantic information.
Figure 2: Illustration of factors influencing the relative importance between short-range and long-range dependencies. The left side presents how the quantitative results vary with changes in the characteristics of the hypothesized factors. The horizontal axis represents the Euclidean distance between the current token and all other tokens, i.e., the range of dependency, while the vertical axis indicates the average importance of other dependencies located at the corresponding distance to the current token. The curve is derived from all tokens across all images in the test set. The right side visualizes the importance of other tokens to the chosen tokens (marked with pentagrams). In (a), the tokens are located at the same position and correspond to the same ground truth image but exhibit different haze densities. In (b), the tokens differ in semantic content but are located within the same image under homogeneous haze conditions.
Method
Figure 3: (a) The overall architecture of DehazeMatic. (b) The internal structure of the Transformer-Mamba Dual Aggregation (TrambaDA) block. (c) The process by which the CLIP-enhanced Dual-path Aggregator (CedA) generates the aggregation weights used to fuse the dual-path outputs of the TrambaDA block. To improve computational efficiency, the TrambaDA blocks share the same aggregation weights within each stage.
We propose DehazeMatic, as illustrated in Figure 3. It adopts a symmetric encoder-decoder framework in which each stage is constructed using a dual-stream design. To effectively aggregate the outputs from the two streams, we introduce the CLIP-enhanced Dual-path Aggregator, which utilizes the estimated haze density and semantic maps to instruct the aggregation process.
Transformer-Mamba Dual Aggregation Block
@@
The TrambaDA block adopts a dual-stream design to simultaneously capture dependencies at different ranges, aiming to provide complementary restoration information for image tokens with distinct characteristics.
The specific feature extractors for each path are flexible and not the focus of this paper. Therefore, for the sake of design simplicity, we adopt window-based self-attention  (Liu et al. 2021) to capture short-range dependencies, with its output denoted as $F_{\mathrm{out\_{short}}}$. To model long-range dependencies, we apply Mamba’s S6 block  (Gu and Dao 2023), equipped with a four-directional scanning strategy (Liu et al. 2024), and denote its output as $F_{\mathrm{out\_{long}}}$. The detailed structure is illustrated in Figure 3(b).
CLIP-enhanced Dual-path Aggregator
Based on the aforementioned insights, CedA is designed to first generate fine-grained haze density and semantic maps, and then produce pixel-level weights accordingly to adaptively aggregate the dual-path outputs, i.e.,
@@
$W_{short},W_{long}\in\mathbb{R}^{H\times W}$ are the aggregation weights corresponding to short- and long-range dependencies, respectively, and their generation process is shown in Figure 3(c).
Estimating Haze Density Map
We begin by defining the estimation pipeline. Inspired by Figure 1, we leverage CLIP to estimate the haze density map. Specifically, given a hazy image $I_{\mathrm{haze}}\in\mathbb{R}^{H\times W\times 3}$ and a paired prompts describing haze and clear conditions $T=[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ (e.g., [’hazy image’, ’clear image’]), we encode them into latent space using a pretrained CLIP model:
@@
To optimize estimation performance and reduce reliance on time-consuming prompt engineering, we employ learnable prompt tokens rather than manually predefined prompts to characterize abstract haze and clear conditions.
The training comprises two stages. In Stage 1, we optimize learnable paired prompts using cross-entropy loss, enabling them to preliminarily distinguish between hazy and clear images. Given a hazy and clear image $I_{\mathrm{haze}},I_{\mathrm{clear}}\in\mathbb{R}^{H\times W\times 3}$, the first-stage loss $\mathcal{L}_{1}$ is defined as:
@@
In Stage 2, the objective is to predict haze density more accurately. The most straightforward and effective optimization approach is regression. However, existing datasets lack ground-truth density maps corresponding to hazy images. To address this, we construct triplet data $\{I_{\mathrm{haze}},I_{\mathrm{clear}},I_{\mathrm{density}}\}$ based on Atmospheric Scattering Model (Narasimhan and Nayar 2002). The second-stage loss $\mathcal{L}_{2}$ is then defined as: